# Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataξ,γ          Muhammad Abdul-Mageedλ,ξ,γ

### v1

@@ 5 Conclusion
We introduced a novel methodology for adapting LLMs to specific communities by deeply integrating their unique linguistic characteristics, cultural heritage, and societal values. Our approach leverages a teacher model proficient in generating low-resource languages to enable: (i) translation for the incorporation of community-specific language and (ii) controlled generation and retrieval mechanisms for the authentic inclusion of cultural heritage and values. We validated our methodology using the Moroccan and Egyptian Arabic dialects as testbeds by developing NileChat, an LLM covering these two dialects. Comprehensive evaluations on understanding, translation, and cultural alignment benchmarks demonstrate that our method significantly enhances the baseline LLM’s performance in capturing target language nuances and cultural values. Notably, NileChat also outperform existing Arabic-aware LLMs. Our method offers a promising research direction for fostering inclusivity of diverse local communities within LLM development, thereby emphasizing the critical role of such an inclusion in the broader democratization of this technology.
@@
  • Computational Cost of Synthetic Data Generation: The process of generating synthetic data is computationally intensive, particularly when employing large teacher models (e.g., Command-R+, a 104B parameter model requiring substantial GPU resources: 4x80GB). This challenge is amplified by the autoregressive generation of long documents from extensive input contexts (e.g., articles, persona descriptions, cultural concepts) restricting the scale of this approach for more languages.
@@ Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^6] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^7] https://arc.ubc.ca/ubc-arc-sockeye
@@ B.1 Pre-training
Details for continued model pretraining. We continue the pre-training of Qwen-2.5-3B on our curated pre-training dataset. Subsequently, the model is fully fine-tuned for one epoch using a sequence length of 4,096. To optimize the learning process, the learning rate is linearly decayed from $5\times 10^{-6}$ to $5\times 10^{-7}$. To mitigate overfitting, we apply a weight decay of $0.1$, and gradient norms are clipped at a maximum value of $1.0$. The training is performed on a cluster of 4$\times$×A100 80GB GPUs.
@@ B.2 Supervised fine-tuning
Table B.1: Comparison of the performance of the Egyptian SFT model (NileChat-EGY), the Moroccan SFT model (NileChat-MOR), and their merged version, NileChat, on Belebele (accuracy), Flores (chrf), and In-house parallel data (chrf).
@@ Appendix D Full results
Less than 7B Qwen3-1.7B 28.53 28.53 28.07 27.33 22.89 22.89
@@
More than 7B AceGPT-7B-chat 40.76 37.98 33.04 31.04 38.00 33.00
@@
Less than 7B Qwen3-1.7B 19.49 19.80 36.32 36.45 14.76 15.45 40.21 40.79 12.17 12.90 35.08 35.83 19.38 19.31 36.23 36.22 19.65 20.48 46.49 47.05 7.32 7.77 22.02 23.42 10.02 10.62 25.98 27.55 5.98 6.63 19.60 21.69 13.59 14.12 32.74 33.62
@@
More than 7B AceGPT-7B-chat 19.92 18.93 36.91 35.72 17.67 20.04 44.09 45.71 14.78 15.92 38.08 38.87 19.11 17.99 36.04 35.24 23.11 25.81 50.28 51.96 8.93 12.29 21.59 29.72 16.11 18.25 33.32 35.56 5.95 10.28 16.65 27.07 15.70 17.44 34.62 37.48
@@
Less than 7B Qwen3-1.7B 12.14 13.14 31.46 31.98 19.11 20.32 34.24 35.09 4.63 4.96 15.10 15.43 2.52 2.64 11.14 10.99 6.19 6.12 22.58 23.01 6.45 6.04 19.25 18.12 18.20 18.26 34.20 33.86 6.20 6.07 18.38 17.94 9.43 9.69 23.29 23.30
@@
More than 7B AceGPT-7B-chat 19.27 21.44 39.36 40.93 20.94 21.99 35.99 37.51 9.24 10.57 22.39 25.25 3.36 3.69 12.76 15.12 7.80 9.95 26.08 27.97 7.15 8.23 20.20 22.17 20.22 20.47 37.14 37.05 6.53 6.00 19.31 18.98 11.81 12.79 26.65 28.12
@@
Table D.2: Detailed translation evaluation results on Flores and In-House datasets. Metrics: spBLEU (sB) and chrF (cF) for 0-shot (0) and 4-shot (4) scenarios. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column.

### v3

@@ 5 Conclusion
We introduced a novel methodology for adapting LLMs to specific communities by deeply integrating their unique linguistic characteristics, cultural heritage, and societal values. Our approach leverages a teacher model proficient in generating low-resource languages to enable: (i) translation for the incorporation of community-specific language and (ii) controlled generation and retrieval mechanisms for the authentic inclusion of cultural heritage and values. We validated our methodology using the Moroccan and Egyptian Arabic dialects as testbeds by developing NileChat, an LLM covering these two dialects. Comprehensive evaluations on understanding, translation, and cultural alignment benchmarks demonstrate that our method significantly enhances the baseline LLM’s performance in capturing target language nuances and cultural values. Notably, NileChat also outperforms existing Arabic-aware LLMs. Our method offers a promising research direction for fostering inclusivity of diverse local communities within LLM development, thereby emphasizing the critical role of such an inclusion in the broader democratization of this technology.
@@
  • Computational Cost of Synthetic Data Generation: The process of generating synthetic data is computationally intensive, particularly when employing large teacher models (e.g., Command R+, a 104B parameter model requiring substantial GPU resources: 4x80GB). This challenge is amplified by the autoregressive generation of long documents from extensive input contexts (e.g., articles, persona descriptions, cultural concepts) restricting the scale of this approach for more languages.
@@ Acknowledgments
Muhammad Abdul-Mageed acknowledges support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^7] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^8] https://arc.ubc.ca/ubc-arc-sockeye
@@ B.1 Pre-training
[…]
@@ B.2 Supervised fine-tuning
Table B.1: Comparison of the performance of the Egyptian SFT model (NileChat-EGY), the Moroccan SFT model (NileChat-MOR), and their merged version, NileChat, on Belebele (accuracy), Flores (ChrF++), and In-house parallel data (ChrF++).
@@ C.1 Evaluation Tasks
The Quality and Validation of Generated Datasets. We rigorously validated the two evaluation Egyptian Arabic datasets we created using machine translation, namely EgyMMLU and EgyHellaSwag. A random sample of items from each dataset was rated on two 1–5 scales: (i) Correctness (semantic accuracy/faithfulness) and (ii) Dialectness (authenticity and naturalness in Egyptian Arabic). Expert human annotator rated 100 randomly sampled items per dataset. To increase statistical power, a state-of-the-art LLM judge (Gemini 2.5 Pro) rated the same 100 items plus an additional 300 (400 total) per dataset. Agreement between human and LLM scores on the 100 shared items yielded an Intraclass Correlation Coefficient (ICC) of 0.60, indicating good reliability. Summary scores appear in Table C.1. These results, together with the inter-evaluator agreement, indicate that the newly created Egyptian benchmarks are of high quality and suitable for reliable model evaluation.
Dataset Evaluator Sample Size Correctness Score Dialectness Score
EgyMMLU Human 100 3.78 3.92
LLM 100 3.90 4.22
LLM 400 4.28 4.22
EgyHellaSwag Human 100 3.90 4.08
LLM 100 4.22 4.96
LLM 400 4.11 4.64
Table C.1: Translation quality scores (out of 5) for EgyMMLU and EgyHellaSwag along two dimensions: correctness and dialectness. Human raters and an LLM judge rated the same 100 items; the LLM judge additionally rated 300 more items (400 total) to yield more reliable estimates.
@@ Appendix D Full results
Less than 7B
Qwen3-1.7B 28.53 28.53 28.07 27.33 22.89 22.89
@@
More than 7B
AceGPT-7B-chat 40.76 37.98 33.04 31.04 38.00 33.00
@@
Less than 7B
Qwen3-1.7B 19.49 19.80 36.32 36.45 14.76 15.45 40.21 40.79 12.17 12.90 35.08 35.83 19.38 19.31 36.23 36.22 19.65 20.48 46.49 47.05 7.32 7.77 22.02 23.42 10.02 10.62 25.98 27.55 5.98 6.63 19.60 21.69 13.59 14.12 32.74 33.62
@@
More than 7B
AceGPT-7B-chat 19.92 18.93 36.91 35.72 17.67 20.04 44.09 45.71 14.78 15.92 38.08 38.87 19.11 17.99 36.04 35.24 23.11 25.81 50.28 51.96 8.93 12.29 21.59 29.72 16.11 18.25 33.32 35.56 5.95 10.28 16.65 27.07 15.70 17.44 34.62 37.48
@@
Less than 7B
Qwen3-1.7B 12.14 13.14 31.46 31.98 19.11 20.32 34.24 35.09 4.63 4.96 15.10 15.43 2.52 2.64 11.14 10.99 6.19 6.12 22.58 23.01 6.45 6.04 19.25 18.12 18.20 18.26 34.20 33.86 6.20 6.07 18.38 17.94 9.43 9.69 23.29 23.30
@@
More than 7B
AceGPT-7B-chat 19.27 21.44 39.36 40.93 20.94 21.99 35.99 37.51 9.24 10.57 22.39 25.25 3.36 3.69 12.76 15.12 7.80 9.95 26.08 27.97 7.15 8.23 20.20 22.17 20.22 20.47 37.14 37.05 6.53 6.00 19.31 18.98 11.81 12.79 26.65 28.12
@@
Table D.2: Detailed translation evaluation results on Flores and In-House datasets. Metrics: spBLEU (sB) and ChrF++ (cF) for 0-shot (0) and 4-shot (4) scenarios. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column.